I am not understanding how to express each power asa radical,then evaluate. (-7776)^2/5
I go over other examples but not understanding how they get the last bit of the answer on the calculator.
They have as an example..... (-243)^1/5=5√-243 = -3
not sure how they got =-3? very stumped....

Answer by Alan3354(69443)

(-3)^5 = -243
So the 5th root of -243 is -3.
-------------------
6^5 = 7776
Odd powers of negatives are negative.
(-6)^5 = -7776
---------------
(-7776)^(1/5) = -6
(-6)^2 = +36
---------------
Calculators use logs for internal calculation, which don't handle negative numbers in most cases (most calculators).